A 0.3-kg block is moving to the right on a horizontal, frictionless surface with a speed of 0.6 m/s. It makes a head-on collision with a 0.2-kg block that is moving to the left with a speed of 1.2 m/s. Find the final velocity (magnitude and direction) of each block if the collision is elastic. (Since the collision is head-on, all motion is along a line.)


1
Expert's answer
2021-05-14T09:42:58-0400

For an elastic collision


"v_1=\frac{m_1-m_2}{m_1+m_2}\cdot u_1+\frac{2m_2}{m_1+m_2}\cdot u_2="


"=\frac{0.3-0.2}{0.3+0.2}\cdot 0.6+\frac{2\cdot0.2}{0.3+0.2}\cdot (-1.2)=-0.84\ (m/s)" to the left


"v_2=\frac{m_2-m_1}{m_1+m_2}\cdot u_2+\frac{2m_1}{m_1+m_2}\cdot u_1="


"=\frac{0.2-0.3}{0.3+0.2}\cdot (-1.2)+\frac{2\cdot0.3}{0.3+0.2}\cdot 0.6=0.96\ (m/s)" to the right